Two books cost 220 Philippine peso and one cost 20 Philippine peso more than the other. Find the cost of each.
There are 10 coins in a purse consisting of 5 Philippine peso and 1 Philippine peso denominations. If the total amount of the coins is 26 Philippine peso then how many coins of each kind are there in the purse.
Thank you so much for the help. This will help me in my entrepreneurship strand.

Answer by ikleyn(52775)

.
1. Two books cost 220 Philippine peso and one cost 20 Philippine peso more than the other. Find the cost of each.

x + y = 220,    (1)
x - y =  20.    (2)

To solve, add equations (1) and (2).

Then, I think, everybody in Philippine knows what to do next.


2. There are 10 coins in a purse consisting of 5 Philippine peso and 1 Philippine peso denominations.
If the total amount of the coins is 26 Philippine peso then how many coins of each kind are there in the purse.

 x +  y = 10,     (1)    (there are 10 coins, x of 5-peso and y of 1-peso)
5x + 1y = 26.     (2)    (the "value" equation)

Distract equation (1) from equation (2).

From this point, complete the assignment yourself.  Good luck!